OpenStudy (anonymous):

PLEASE HELP WITH THIS CALCULUS PROBLEM! A carousel has a diameter of 16m and completes a revolution every 30s. a) Model the north-south position of a rider on the outside rim of the carousel using a sine function. b) Differentiate the function in part a) c) Determine the north-south speed of the rider. d) What is the position of the rider on the carousel when this maximum north-south speed is reached? PLEASE SHOW WORK!

OpenStudy (anonymous):

a) This is uniform circular motion but we can model the x motion and the y motion separately. Calling y>0 north and y<0 south, \[y=8\sin wt\]where t=time in seconds and w=angular velocity in radians/second. The 8 appears in the question because the north-south position varies between +8m and -8m since the diameter of the wheel is 16m. The angular velocity is given by:\[w=\frac{2pi}{30}=\frac{pi}{15}\]so we have:\[y=8\sin (\frac{pi}{15}t)\]

OpenStudy (anonymous):

Ok, thank you for that, now to part b

OpenStudy (anonymous):

give part b a try and let me know if you get stuck

OpenStudy (anonymous):

I dont know how to differentiate it

OpenStudy (anonymous):

do you know how to differentiate \[y=\sin x\]?

OpenStudy (anonymous):

cosx

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

now y=8sinx just differentiates to 8cosx so we're almost there.

OpenStudy (anonymous):

so its just 8cosx(pi/15)(t)?

OpenStudy (anonymous):

not quite; we need to apply the chain rule to this one

OpenStudy (anonymous):

\[\frac{d}{dt}8\sin (\frac{pi}{15}t)=8\cos(\frac{pi}{15}t)*\frac{d}{dt}(\frac{pi}{15}t)\]

OpenStudy (anonymous):

so is it: 8/15picos(pi(t)/15)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

\[dy/dt=\frac{8pi}{15}\cos (\frac{pi}{15}t)\]yeah

OpenStudy (anonymous):

so how about part c)?

OpenStudy (anonymous):

Well part c is badly worded. Do they want the value of the maximum speed? The north-south speed is given by \[\left| dy/dt \right|\] which varies with time. So the north-south speed is just the absolute value of the function in part b.

OpenStudy (anonymous):

im not sure, thats how its written

OpenStudy (anonymous):

i think its asking for maximum according to part d)

OpenStudy (anonymous):

my answer would be \[\left| \frac{8pi}{15}\cos (\frac{pi}{15}t) \right|\]Now the largest the absolute value of the cosine of anything can be is 1...this occurs when t=0 and t=-pi and t=2pi. This max speed is equal to (8*pi)(15)

OpenStudy (anonymous):

are you sure thats the answer?

OpenStudy (anonymous):

If you see an error, let me know

OpenStudy (anonymous):

ok, onto part d)

OpenStudy (anonymous):

which equation gives us the position of the rider?

OpenStudy (anonymous):

Im not sure

OpenStudy (anonymous):

btw, I got 377 for (8*pi)(15). Whats the unit?

OpenStudy (anonymous):

ok, we already came up with it...it is precisely the equation asked for in part a. Now, we know the values of t for which the north-south velocity is at a maximum so all you need to do is plug in these value of t into the position function to answer this.

OpenStudy (anonymous):

units are in m/s

OpenStudy (anonymous):

I dont understand what values of t you are talking about

OpenStudy (anonymous):

and (8*pi)/15 is about 1.68...not sure where you got 377?

OpenStudy (anonymous):

"This max speed is equal to (8*pi)(15)"

OpenStudy (anonymous):

oh, that's a typo on my part...should read (8*pi)/15

OpenStudy (anonymous):

ok so what values of t?

OpenStudy (anonymous):

"Now the largest the absolute value of the cosine of anything can be is 1...this occurs when t=0 and t=-pi and t=2pi. This max speed is equal to (8*pi)(15) "

OpenStudy (anonymous):

so I just substitute all those values and thats the answer?

OpenStudy (anonymous):

substitute them into the equation from part a, and yes, those should be the answers

OpenStudy (anonymous):

when writing pi, is it 3.14 or 180?

OpenStudy (anonymous):

3.14 radians

OpenStudy (anonymous):

so when t = 0, position = 0?

OpenStudy (anonymous):

the y position equals zero, yep

OpenStudy (anonymous):

and when t= -pi, y= -0.9186?

OpenStudy (anonymous):

no...when t=-pi, y=0 again

OpenStudy (anonymous):

how?

OpenStudy (anonymous):